# Staying on Top of SMEFT-Likelihood Analyses

### v1

@@ Abstract
We present a new global SMEFT analysis of LHC data in the top sector. After updating our set of measurements, we show how public ATLAS likelihoods can be incorporated in an external global analysis and how our analysis benefits from the additional information. We find that, unlike for the Higgs and electroweak sector, the SMEFT analysis of the top sector is mostly limited by the correlated theory uncertainties.
@@ 1 Introduction
Over the last decade, LHC physics has seen a paradigm shift, from testing models for physics beyond the Standard Model (BSM) to precision measurements and completely understanding LHC physics in terms of fundamental quantum field theory. This change represents impressive progress in experiment and theory, in a holistic approach to the huge LHC dataset. On the theory side, a driving force is the development of an effective field theory version of the Standard Model (SMEFT), which allows us to ask, and answer, the question: Does the LHC data agree with the Standard Model altogether?
SMEFT is a perturbative quantum field theory that respects the gauge symmetries and covers all sectors of the Standard Model. It is renormalizable and allows for QCD and electroweak precision predictions. It is built on the idea that BSM particles affecting LHC measurements might be too heavy to be produced on-shell. Assuming that the Higgs and Goldstone fields form the usual doublet, SMEFT embeds the Standard Model in an effective field theory (EFT). The SMEFT idea [1, 2, 3, 4] was developed for a gauge-invariant description of anomalous gauge interactions at LEP [5, 6]. Its big success is the unified global analysis of the Higgs and electroweak sector, including electroweak precision data [7, 8, 9], a major part of the legacy of the LHC Run 2 [10, 11, 12, 13, 14]. Obviously, the same approach [15, 16, 17] can be used to systematically test the top quark sector [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. This sector is especially interesting, because it can be combined with the bottom sector [29, 30, 31, 32, 33, 34] and a much broader set of precision measurements [35], eventually testing the impact of flavor symmetries. Recently, several groups have provided combined SMEFT analyses of the electroweak and top sectors [36, 37], SMEFT analyses combined with parton density extraction [38, 39, 40, 41], and even SMEFT analyses with lighter new particles [42, 43, 44].
@@
The publication of experimental data in the format of public likelihoods is a major step in the way experimental results can be re-interpreted [50, 51, 52, 53, 54, 55]. The HistFactory format [56], software such as pyhf [57, 58], Spey [59], and simplified likelihood frameworks [60] allow for an efficient use of likelihoods. In the classic BSM sector many likelihoods have already been made public and analyzed [61]. In the top sector, public ATLAS likelihoods [47, 48, 49] still have to be used outside the collaborations. We aim to fill this gap by using them as the basis of a global SMEFT analysis of the top sector using SFitter.
In this paper we update an earlier NLO-SMEFT analysis of the top sector in the SFitter framework [22]. SFitter is unique in that it does not rely on pre-processed experimental measurements and includes its own comprehensive uncertainty treatment [62, 63, 64]. This makes it a promising candidate to, for the first time, include public likelihoods in a global analysis and determine their impact. After a general introduction to the SFitter methodology and our dataset in Sec. 2, we will discuss three public likelihoods in the top sector in detail in Sec. 3.
In our global SFitter analysis we will focus on systematic and theory uncertainties. In addition, we will probe the impact of a profile likelihood vs Bayesian marginalization when extracting limits on single Wilson coefficients, where we saw significant effects on the Higgs and electroweak sector [65]. Finally, we will provide a short comparison between the Markov chains used in this analysis and the Monte Carlo experiment method used in earlier SFitter analyses in the Appendix.
@@ 2.1 SMEFT Lagrangian
By fundamental theory arguments, the SMEFT Lagrangian is the appropriate interpretation framework to interpret LHC searches for effects of particles which are too heavy to be produced on-shell [4]. While in the Higgs sector one can argue about the proper way to implement electroweak symmetry breaking and the doublet nature of the Higgs and Goldstone fields, the SMEFT description of the top sector is fixed. An open question is how to combine it with the light-flavor sector and its range of potential global symmetries. This renders the impact of flavor measurements on the top sector somewhat unclear, so we will not exploit this link and instead refer to dedicated analyses [29, 30, 31, 32, 33].
@@
$\mathscr{L}_{\text{eff}}=\sum_{k}\left(\frac{C_{k}}{\Lambda^{2}}\,^{\ddagger}\mathcal{O}_{k}+\text{h.c.}\right)+\sum_{k}\frac{C_{k}}{\Lambda^{2}}\,\mathcal{O}_{k}\;.$ (1)
This means the sum runs over all operators at mass dimension six, involving top-quarks. Non-hermitian operators are denoted as ${}^{\ddagger}\mathcal{O}$. We neglect operators of mass dimension seven and higher in the EFT expansion, assuming that their $\Lambda$-suppression translates into a suppression of their effects on LHC observables. This assumption is formally well motivated, but given the rather modest scale separation between the LHC and the accessible $\Lambda$-values, it has to be checked for a given dataset and a given UV-completion matched to the SMEFT Lagrangian [67, 68, 69, 70, 71].
Because the underlying symmetry structure is an input to an EFT construction and we are hesitant to leave the test of fundamental symmetries to a numerically tricky and hardly conclusive global analysis [72], we ignore CP-violating operators. Finally, the fact that the top-sector measurements included in our analysis are blind to the light-quark flavor we assume separate $U(2)$ symmetries in the first and second generation [73, 74],
@@
Six operators include a $LR$ or $RL$ chirality flip in the current-current interaction,
@@
Finally, there are eight operators which couple two heavy quarks to the gauge bosons [75],
@@
$C_{\phi Q}^{\mp}=C_{\phi Q}^{1}\mp C_{\phi Q}^{3}\qquad\text{and}\qquad C_{tZ}=c_{w}C_{tW}-s_{w}C_{tB}.$ (6)
This way, $C_{\phi Q}^{-}$ and $C_{tZ}$ describe a $t\bar{t}Z$ interaction, $C_{tW}$ a $tbW$ interaction, and $C_{\phi Q}^{3}$ both $tbW$ and $b\bar{b}Z$ interactions. The effect of our operators on the different LHC observables are summarized in Tab. 1. Here the main question is which operators modify the LHC rate and kinematic predictions through interference with the SM-matrix element and which only contribute at dimension-6 squared.
@@ 2.2 Data, predictions, and uncertainties
[…]
@@
[…]
@@
With this in mind, our analysis in the top sector will consider all measurements listed in Tables 2 and 3. The analysis is an update to a previous global top analysis in SFitter, Ref. [22]. We highlight in Tables 2 and 3 the measurements which are new relative to those included in Ref. [22], as well as those for which a public likelihood is available. Where possible, we make use of measurements encompassing the full Run II LHC luminosity, and choose measurements in the boosted regime in which sensitivity to energy-growing SMEFT operators is maximized; see, for example, the top pair production invariant mass distribution of Ref. [94]. The dataset consists of a total of 122 datapoints spanning the $t\bar{t}$, $t\bar{t}+X(Z,W,\gamma)$ and single top ($s$, $t$-channel, $tW$ and $tZ$) sectors, including measurements of top pair production charge asymmetries $A_{C}$ and $W$ boson polarization in top decays ($F_{0},F_{L}$).
@@
A key ingredient to all global analyses are precision predictions from perturbative quantum field theory. Most observables considered in this analysis are unfolded to parton level, assuming stable top quarks. This allows us to use fixed-order calculations to determine the SM predictions at NLO in QCD using MadGraph5 aMC@NLO [132, 133] and NNPDF 4.0 [134] interfaced with LHAPDF [135]. Alongside the observables listed in Tables 2 and 3, we note whether the SM predictions for these observables is approximated at NNLO in QCD using a $K$-factor approximation, and reference the source of these QCD $K$-factors. In the case of new top quark pair production observables these QCD $K$-factors are calculated using HighTea [45].
@@
Calculations of the effect of the SMEFT on all updated measurements in the top sector are performed at NLO in QCD using the FeynRules [136] model SMEFTatNLO [137] up to quadratic order in the EFT expansion. The exception are the measurements of the $t\bar{t}\gamma$ total cross sections at 8 TeV by ATLAS [107] and CMS [106], for which the SMEFT predictions at LO in QCD are taken from SMEFiT [138, 37].
@@
Theory uncertainties appear whenever we compare a measurement to a first-principle description. In principle, they cover a wide range of approximations which we make to be able to calculate, for example, an LHC cross section from a fundamental renormalized Lagrangian. For the LHC, they are dominated by the truncation of the perturbative series, in QCD and in the electroweak gauge coupling. Because these perturbative series converge very slowly for LHC rates, theory predictions have become limiting factors for the interpretation of many LHC measurements in terms of actual physics. Aside from the size of the theory uncertainties, it is problematic that they do not follow any statistical pattern or model [139], and assuming a Gaussian distribution is neither justified nor conservative.
@@
Because of their impact on global analyses of effective Lagrangians, SFitter puts an emphasis on the proper description of these uncertainties, including their correlations between different observables. We will describe this treatment in more detail in Sec. 2.3. In the top sector the theory uncertainties are critical for the precisely measured top pair production rates [22] and are correlated between different final states for rate measurements. We typically use the theory uncertainties reported in the respective publications, with the exception that we enforce a minimum scale uncertainty of 10% for total rates and 2% for bin-wise kinematic distributions.
@@
allowing for an easy comparison with fixed-order calculations. This, alongside the high-$m_{t\bar{t}}$ reach of this distribution, makes it an excellent candidate for constraining the energy- growing SMEFT four-fermion operators of the top sector. We display the impact of one of these operators, $\mathcal{O}_{Qd}^{8}$, in Fig.1.
The theory uncertainty is shown in blue in both figures, and compared to the statistical and systematic uncertainties on the experimental data. In both cases, the values of $\mathcal{C}_{Qd}^{8}$ chosen are those which would produce a $3\sigma$ deviation in a one-parameter analysis. We observe that, while the measurement unfolded to the full phase space is sensitive to the energy-growing effects of $\mathcal{O}_{Qd}^{8}$, this sensitivity is significantly enhanced by the measurement of boosted top quarks.
@@ 2.3 SFitter
The SFitter framework [62, 63, 64] has been developed for global analyses of LHC measurements in the context of BSM physics and Higgs properties [140, 7, 12, 22], including comprehensive studies on Higgs and electroweak properties induced by actual UV-completion of the SM [141, 69], an extrapolation for the HL-LHC [65], as well as for future electron-positron colliders [142].
The relation to full models and a proper treatment of the uncertainties in precision matching is crucial for the LHC, because the typical scale separation between directly probed energies and indirectly accessible energies is not very large. On the other hand, the consistency of the EFT description is not a universal property of the EFT Lagrangian, but only defined by possible on-shell propagators in the observables and relative to the UV-completion and its typical coupling strengths. Without additional information on the underlying model the Lagrangian in Eq.(1) is degenerate along $C_{k}\sim\Lambda^{2}$, which means the EFT assumption of large $\Lambda$ improves for larger postulated couplings. This is the reason why SFitter SMEFT analyses start with the truncated dimension-6 Lagrangian at face value.
From Tab. 1 we know that some Wilson coefficients do not interfere with the SM matrix elements at leading order, so we include dimension-6 squared contributions to the LHC observables. This means we truncate the Lagrangian rather than the LHC rate prediction. We emphasize that all our assumptions are just that, there is no right or wrong, and they need to be checked for a given dataset and a given UV-completion [141, 67, 68, 69]. However, our assumptions ensure that the SFitter analysis makes optimal use of the kinematic information especially in the tails of momentum or energy distributions.
@@
It describes the statistical, systematic, and theory uncertainties. The first Poisson distribution gives the probability to observe $d$ events given the corresponding theory prediction $m(c,\theta,b)$, which in turn depends on the predicted background count $b$. The background rate is, itself, constrained by measurements $b_{\text{CR}}$ in the control region, implemented as a scaled prediction $kb$ with a suitable factor $k$. The constraint function $\mathcal{C}$ gives the distribution of the nuisance parameter $\theta_{i}$, given a width measure $\sigma_{i}$. Depending on the source of the uncertainty it can be chosen as follows:
@@
The flat scale uncertainty is not parametrization invariant, as one would expect from a fixed range, but without a preferred central value we consider it conservative.
@@
By ansatz, SFitter describes all measurements $d$ as uncorrelated. This is justified by the individual statistical uncertainties described by the Poisson distributions in the likelihood. Correlations can be included through the $\mathcal{C}$ distributions. In the case of Gaussian systematics, the correlations are described by a correlation matrix [143], which we choose as either uncorrelated or (essentially) fully correlated. The implementation of flat theory uncertainties is through an allowed shift of the prediction at no cost in the likelihoods. Theory uncertainties are correlated for all measurements with identical predictions. This is done by averaging them, weighted such that the final standard deviation is minimized. Using the prediction only once for this weighted average, instead of each individual measurement, ensures the proper correlations of the corresponding theory uncertainties.
@@
To construct the exclusive likelihood, SFitter uses cross section predictions over the entire model parameter space and extracts the quadratic behavior analytically, which guarantees sufficient precision even for small Wilson coefficients. We then use a Markov chain to evaluate Eq.(8) numerically, and to encode the likelihood in the distribution of points covering the combined space of Wilson coefficients and nuisance parameters. The setup of the Markov chain depends on the structure of the model space; for BSM analyses an efficient search for local structures in the global model space is important, while for SMEFT analyses we know that the global likelihood maximum will be close to the SM-point. Adjusting the Markov chain accordingly leads to a significant speed improvement [69]. Our motivation to use a Markov chain rather than so-called toys is described in the Appendix.
@@
Finally, to combine uncertainties by removing nuisance parameters, or to reduce the space of physical Wilson coefficients, SFitter can employ a profile likelihood or a Bayesian marginalization [63, 65]. Obviously, these two methods give different results. Only for uncorrelated Gaussians do the profile likelihood and Bayesian marginalization lead to the common result of errors added in quadrature. For a flat likelihood the uncorrelated profile likelihood adds the two uncertainties linearly, which happens for the scale uncertainty and the PDF uncertainty in SFitter. The profile likelihood combination of a flat and a Gaussian uncertainty gives the well-known RFit prescription [144].

### v3

@@ Abstract
We present a new global SMEFT analysis of LHC data in the top sector. After updating our set of measurements, we show how public ATLAS likelihoods can be incorporated into an external global analysis and how our analysis benefits from the additional information. We find that, unlike for the Higgs and electroweak sector, the SMEFT analysis of the top sector is mostly limited by the theory uncertainties. Finally, we present the first global SFitter analysis combining the top and electroweak-Higgs sectors.
@@ 1 Introduction
Over the last decade, LHC physics has seen a paradigm shift, from testing models for physics beyond the Standard Model (BSM) to precision measurements and a complete understanding of LHC physics in terms of fundamental quantum field theory. This change represents impressive progress in experiment and theory in a holistic approach to the huge LHC dataset. On the theory side, a driving force is the development of an effective field theory version of the Standard Model (SMEFT), which allows us to ask and answer the question: Does the LHC data agree with the Standard Model altogether?
SMEFT is a perturbative quantum field theory that respects the gauge symmetries and covers all sectors of the Standard Model. It is renormalizable and allows for QCD and electroweak precision predictions. It is built on the idea that BSM particles affecting LHC measurements might be too heavy to be produced on-shell. Assuming that the Higgs and Goldstone fields form the usual doublet, SMEFT embeds the Standard Model in an effective field theory (EFT). The SMEFT idea [1, 2, 3, 4] was developed for a gauge-invariant description of anomalous gauge interactions at LEP [5, 6]. Its big success is the unified global analysis of the Higgs and electroweak sector, including electroweak precision data [7, 8, 9], a major part of the legacy of the LHC Run 2 [10, 11, 12, 13, 14]. Obviously, the same approach [15, 16, 17] can be used to systematically test the top quark sector [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. This sector is especially interesting because it can be combined with the bottom sector [29, 30, 31, 32, 33, 34] and a much broader set of precision measurements [35], eventually testing the impact of flavor symmetries. Recently, several groups have provided combined SMEFT analyses of the electroweak and top sectors [36, 37], SMEFT analyses combined with parton density extraction [38, 39, 40, 41], and even SMEFT analyses with lighter new particles [42, 43, 44].
@@
The publication of experimental data in the format of public likelihoods is a major step in the way experimental results can be re-interpreted [50, 51, 52, 53, 54, 55]. The HistFactory format [56], software such as pyhf [57, 58], Spey [59], and simplified likelihood frameworks [60] allow for an efficient use of likelihoods. In the classic BSM sector, many likelihoods have already been made public and analyzed [61]. In the top sector, public ATLAS likelihoods [47, 48, 49] still have to be used outside the collaborations. We aim to fill this gap by using them as the basis of a global SMEFT analysis of the top sector using SFitter.
In this paper, we update an earlier NLO-SMEFT analysis of the top sector in the SFitter framework [22]. SFitter is unique in the sense, that it does not rely on pre-processed experimental measurements and includes its own comprehensive uncertainty treatment [62, 63, 64]. This makes it a promising candidate to, for the first time, include public likelihoods in a global analysis and determine their impact. After a general introduction to the SFitter methodology and our dataset in Sec. 2, we will discuss three public likelihoods in the top sector in detail in Sec. 3. We will then include these likelihoods in the first SFitter analysis of the electroweak and top sectors in Sec. 4. While the physics behind combining these two sectors is largely understood [36, 37], in our global SFitter analysis, we will focus on the impact of theory uncertainties. In addition, we will probe the impact of a profile likelihood vs Bayesian marginalization when extracting limits on single Wilson coefficients, where we saw significant effects on the Higgs and electroweak sector [65]. Finally, we will provide a short comparison between the Markov chains used in this analysis and the Monte Carlo experiment method used in earlier SFitter analyses in the Appendix.
@@ 2.1 SMEFT Lagrangian
By fundamental theory arguments, the SMEFT Lagrangian is the appropriate interpretation framework to interpret LHC searches for effects of particles which are too heavy to be produced on-shell [4]. While in the Higgs sector one can argue about the proper way to implement electroweak symmetry-breaking and the doublet nature of the Higgs and Goldstone fields, the SMEFT description of the top sector is fixed. An open question is how to combine it with the light-flavor sector and its range of potential global symmetries. This renders the impact of flavor measurements on the top sector somewhat unclear, so we will not exploit this link and instead refer to dedicated analyses [29, 30, 31, 32, 33].
@@
$\mathscr{L}_{\text{eff}}=\sum_{j}\left(\frac{C_{j}}{\Lambda^{2}}\,^{\ddagger}\mathcal{O}_{j}+\text{h.c.}\right)+\sum_{k}\frac{C_{k}}{\Lambda^{2}}\,\mathcal{O}_{k}\;.$ (1)
This means the sum runs over all operators at mass dimension six, involving top quarks. Non-hermitian operators are denoted as ${}^{\ddagger}\mathcal{O}$. We neglect the Weinberg operator at dimension 5, as well as all operators of mass dimension seven and higher in the EFT expansion, assuming that their $\Lambda$-suppression translates into a suppression of their effects on LHC observables. This assumption is formally well-motivated but given the rather modest scale separation between the LHC and the accessible $\Lambda$-values, it has to be checked for a given dataset and a given UV-completion matched to the SMEFT Lagrangian [67, 68, 69, 70, 71].
Because the underlying symmetry structure is an input to an EFT construction, and we are hesitant to leave the test of fundamental symmetries to a numerically tricky and hardly conclusive global analysis [72], we ignore CP-violating operators. Finally, the fact that the top-sector measurements included in our analysis are blind to the light-quark flavor we assume separate $U(2)$ symmetries in the first and second generation [73, 74],
@@
Six operators show a $LR$ or $RL$ chirality in the current-current interaction,
@@
Finally, there are eight operators, which couple two heavy quarks to the gauge bosons [75],
@@
$C_{\phi Q}^{\pm}=C_{\phi Q}^{1}\pm C_{\phi Q}^{3}\qquad\text{and}\qquad C_{tZ}=c_{w}C_{tW}-s_{w}C_{tB}.$ (6)
This way, $C_{\phi Q}^{-}$ and $C_{tZ}$ describe a $t\bar{t}Z$ interaction, $C_{tW}$ a $tbW$ interaction, and $C_{\phi Q}^{3}$ both $tbW$ and $b\bar{b}Z$ interactions. The effect of our operators on the different LHC observables are summarized in Tab. 1. Here the main question is which operators modify the LHC rate and kinematic predictions through interference with the SM-matrix element which only contributes at dimension-6 squared order.
@@ 2.2 Data, predictions, and uncertainties
Experiment Energy [TeV] $\mathcal{L}$ [fb-1] Channel Observable # Bins New Likelihood QCD k-factor
@@
Exp. $\sqrt{s}$ [TeV] $\mathcal{L}$ [fb-1] Channel Observable # Bins New Likelihood QCD k-factor
@@
With this in mind, our analysis in the top sector will consider all measurements listed in Tables 2 and 3. The analysis is an update to a previous global top analysis performed by SFitter, in Ref. [22]. We highlight in Tables 2 and 3 the measurements that are new relative to those included in Ref. [22], as well as those for which a public likelihood is available. Where possible, we make use of measurements encompassing the full Run II LHC luminosity and choose measurements in the boosted regime in which sensitivity to energy-growing SMEFT operators is maximized; see, for example, the top pair production invariant mass distribution of Ref. [94]. The dataset consists of a total of 122 data points spanning the $t\bar{t}$, $t\bar{t}+X(Z,W,\gamma)$ and single top ($s$, $t$-channel, $tW$ and $tZ$) sectors, including measurements of top-pair production charge asymmetries $A_{C}$ and $W$ boson polarization in top decays ($F_{0},F_{L}$).
@@
A key ingredient to all global analyses are precision predictions from perturbative quantum field theory. Most observables considered in this analysis are unfolded to parton level, assuming stable top quarks. This allows us to use fixed-order calculations to determine the SM predictions at NLO in QCD using MadGraph5 aMC@NLO [132, 133] and NNPDF 4.0 [134] interfaced with LHAPDF [135]. Alongside the observables listed in Tables 2 and 3, we note whether the SM predictions for these observables are approximated at NNLO in QCD using a $K$-factor approximation and referencing the source of these QCD $K$-factors. In the case of new top quark pair production observables these QCD $K$-factors are calculated using HighTea [45].
@@
Calculations of the effect of the SMEFT on all updated measurements in the top sector are performed at NLO in QCD using the FeynRules [136] model SMEFTatNLO [137] up to quadratic order in the EFT expansion. The exceptions are the measurements of the $t\bar{t}\gamma$ total cross sections at 8 TeV by ATLAS [107] and CMS [106], for which the SMEFT predictions at LO in QCD are taken from Ref. [39].
@@
Theory uncertainties appear whenever we compare a measurement to a first-principle description. In principle, they cover a wide range of approximations which we make to be able to calculate, for example, an LHC cross section from a fundamental renormalized Lagrangian. For the LHC, they are dominated by the truncation of the perturbative series, in QCD and the electroweak gauge coupling. Because these perturbative series converge very slowly for LHC rates, theory predictions have become limiting factors for the interpretation of many LHC measurements in terms of actual physics. Aside from the size of the theory uncertainties, it is problematic that they do not follow any statistical pattern or model [138], and assuming a Gaussian distribution is neither justified nor conservative.
@@
Because of their impact on global analyses of effective Lagrangians, SFitter puts an emphasis on the proper description of these uncertainties, including their correlations between different observables. We will describe this treatment in more detail in Sec. 2.3. In the top sector, the theory uncertainties are critical for the precisely measured top pair production rates [22] and are correlated between different final states for rate measurements. We typically use the theory uncertainties reported in the respective publications, with the exception that we enforce a minimum scale uncertainty of 10% for total rates in single top production and 2% for bin-wise kinematic distributions.
@@
allowing for an easy comparison with fixed-order calculations. This, alongside the high-$m_{t\bar{t}}$ reach of this distribution makes it an excellent candidate for constraining the energy- growing SMEFT four-fermion operators of the top sector. We display the impact of one of these operators, $\mathcal{O}_{Qd}^{8}$, in Fig.1.
The theory uncertainty is shown in blue in both figures and compared to the statistical and systematic uncertainties in the experimental data. In both cases, the values of $\mathcal{C}_{Qd}^{8}$ chosen are those which would produce a $3\sigma$ deviation in a one-parameter analysis. We observe that, while the measurement unfolded to the full phase space is sensitive to the energy-growing effects of $\mathcal{O}_{Qd}^{8}$, this sensitivity is significantly enhanced by the measurement of boosted top quarks.
@@ 2.3 SFitter
The SFitter framework [62, 63, 64] has been developed for global analyses of LHC measurements in the context of BSM physics and Higgs properties [139, 140, 7, 12, 22], including comprehensive studies on Higgs and electroweak properties induced by actual UV-completion of the SM [141, 69], an extrapolation for the HL-LHC [65], as well as for future electron-positron colliders [142].
The relation to full models and a proper treatment of uncertainties in precision matching is crucial for the LHC because the typical scale separation between directly probed energies and indirectly accessible energies is not very large. On the other hand, the consistency of the EFT description is not a universal property of the EFT Lagrangian, but only defined by possible on-shell propagators in the observables and relative to the UV-completion and its typical coupling strengths. Without additional information on the underlying model the Lagrangian in Eq.(1) is degenerate along $C_{k}\sim\Lambda^{2}$, which means the EFT assumption of large $\Lambda$ improves for larger postulated couplings. This is the reason why SFitter SMEFT analyses start with the truncated dimension-6 Lagrangian at face value.
From Tab. 1, we know that some Wilson coefficients do not interfere with the SM matrix elements at leading order, so we include dimension-6 squared contributions to the LHC observables. This means we truncate the Lagrangian rather than the LHC rate prediction. We emphasize that all our assumptions are neither inherently right nor wrong, and need to be validated for a given dataset and a given UV-completion [141, 67, 68, 69]. However, our assumptions ensure that the SFitter analysis makes optimal use of the kinematic information, especially in the tails of momentum or energy distributions.
@@
It incorporates the effects of the statistical, systematic, and theory uncertainties. The first Poisson distribution gives the probability to observe $d$ events given the corresponding theory prediction $m(c,\theta,b)$, which in turn depends on the predicted background count $b$. The background rate is, itself, constrained by measurements $b_{\text{CR}}$ in the control region, implemented as a scaled prediction $kb$ with a suitable factor $k$. The constraint function $\mathcal{C}$ gives the distribution of the nuisance parameter $\theta_{i}$, given a width measure $\sigma_{i}$. Depending on the source of the uncertainty, it can be chosen as follows:
@@
The flat scale uncertainty is not parametrization invariant, as one would expect from a fixed range, but without a preferred central value we consider it conservative. Scale uncertainties are obtained by varying the renormalization and factorization scales $\mu_{R}$ and $\mu_{F}$ by a factor of 2 around their respective central value. These are process dependent and chosen to be $\mu_{R}=\mu_{F}=m_{t}+\frac{1}{2}m_{V}$ for associated $t\bar{t}$ production with $V=W,Z$. For $t\bar{t}$ production, the sum of the transverse masses of the top and anti-top is used, while for single top production they are set to the top mass $m_{t}$.
By ansatz, SFitter treats all measurements $d$ as uncorrelated, constructing an individual likelihood for each measurement, as defined in Eq.(8). This assumption is justified by the individual statistical uncertainties described by the Poisson distributions in the likelihood. The full likelihood can thus be constructed as the product of these individual contributions. As a consequence, the Gaussian constraint terms describing the systematics of each measurement can be generalized to a single higher-dimensional Gaussian which allows correlations between uncertainties to be introduced. In the case of Gaussian systematics, the correlations are described by a correlation matrix [143]
$C_{ij}=\frac{\sum_{\text{syst}}\rho_{ij}\sigma_{i,\text{syst}}\sigma_{j,\text{syst}}}{\sigma_{i,\text{exp}}\sigma_{j,\text{exp}}}\qquad\text{with}\qquad\sigma_{i,\text{exp}}^{2}=\sum_{\text{syst}}\sigma_{i,\text{syst}}^{2}+\sum_{\text{Poiss}}\sigma_{i,\text{Poiss}}^{2}\,.$ (9)
Here $i,j$ run over all measurements and $\sigma_{\text{syst}},\sigma_{\text{Poiss}}$ are the systematic and Poisson uncertainties. We then choose $\rho_{ij}$ to be either uncorrelated or (essentially) fully correlated for systematics of the same type.
@@
Theory uncertainties are correlated for all measurements with identical predictions. They are also correlated within one measurement across all bins but not across several different measurements. This is done by averaging them, weighted such that the final standard deviation is minimized. Using the prediction only once for this weighted average, instead of each individual measurement, ensures the proper correlations of the corresponding theory uncertainties. The implementation of flat theory uncertainties allows for a shift of the prediction within their bounds without affecting the likelihood value.
@@
To construct the exclusive likelihood, SFitter uses cross section predictions over the entire model parameter space and extracts the quadratic behavior analytically, which guarantees sufficient precision even for small Wilson coefficients. We then use a Markov chain to evaluate Eq.(8) numerically and to encode the likelihood in the distribution of points covering the combined space of Wilson coefficients and nuisance parameters. The setup of the Markov chain depends on the structure of the model space; for BSM analyses an efficient search for local structures in the global model space is important, while for SMEFT analyses we know that the global likelihood maximum will be close to the SM-point. Adjusting the Markov chain accordingly leads to a significant speed improvement [69]. Our motivation to use a Markov chain rather than so-called toys is described in the Appendix.
@@
Finally, to combine uncertainties by removing nuisance parameters or to reduce the space of physical Wilson coefficients, SFitter can employ a profile likelihood or a Bayesian marginalization [63, 65]. Obviously, these two methods give different results. Only for uncorrelated Gaussians do the profile likelihood and Bayesian marginalization lead to the common result of errors added in quadrature. For a flat likelihood, the uncorrelated profile likelihood adds the two uncertainties linearly, which happens for the scale uncertainty and the PDF uncertainty in SFitter. The profile likelihood combination of a flat and a Gaussian uncertainty gives the well-known RFit prescription [144]. In contrast, when applying marginalization on the combination of Gaussian and flat uncertainties, the central limit theorem ensures that the final posterior will be Gaussian again.